OpenStudy (anonymous):

Johns father is head manager of a company that produces a best selling product. During the most recent computer upgrade, a new production inequality,|-2x-4| less than or equal to 12, was added to speed up production. The variable x in the inequality represents the amount of raw material ( in kilograms) to make the product.

OpenStudy (anonymous):

And what does the question ask?

OpenStudy (anonymous):

1.Solve the inequality 2.if x represents positive amounts in kilograms, what is the solution set? 3.only positive integers solutions can be used and accepted. How does this change the answer to #2. Explain

OpenStudy (anonymous):

\[|-2x - 4|\le 12\] \[-12 \le -2x - 4 \le 12\]

OpenStudy (anonymous):

From there you can arrange the terms and solve this inequality.